40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 26 17:24:24 2017, Steve, Update, General, power outage powerOutage.png
    Reply  Thu Dec 28 17:22:02 2017, gautam, Update, General, power outage - CDS recovery 28.png
       Reply  Thu Jan 4 14:39:50 2018, gautam, Update, General, power outage - timing error 
          Reply  Fri Jan 5 21:54:28 2018, rana, Update, General, power outage - timing error huh.png
             Reply  Fri Jan 5 22:19:53 2018, gautam, Update, General, power outage - timing error 
                Reply  Sat Jan 6 18:27:37 2018, gautam, Update, General, power outage - IFO recovery EX_QPD_Quad1_Faulty.pdfTrillium_faulty.pdf
                   Reply  Mon Apr 16 22:09:53 2018, Kevin, Update, General, power outage - BLRM recovery AA.jpgADC.jpg
                      Reply  Tue Apr 17 09:57:09 2018, Steve, Update, General, seismometer interfaces BS_Tril_Intrf-1X5.jpgGurs_Intf-1X1.jpg
                         Reply  Wed Apr 18 20:33:19 2018, Kevin, Update, General, seismometer interfaces front.jpgback.jpgconnectors.jpg
                            Reply  Fri Apr 20 23:36:28 2018, Kevin, Update, PEM, Seismometer BLRMs front.jpgback.jpgdetail.jpg
                               Reply  Tue Apr 24 21:19:08 2018, Kevin, Update, PEM, Seismometer BLRMs modified_schematic.pdftop.jpgbottom.jpgBS_Seis_PSD.pdf
                                  Reply  Wed Apr 25 17:44:39 2018, Arnold, Update, PEM, PEM Anti-Alias wiring 
                                     Reply  Thu Apr 26 09:35:49 2018, Kevin, Update, PEM, PEM Anti-Alias wiring board.jpgSeismometerPSD.pdf
                                        Reply  Thu Apr 26 20:22:21 2018, Kevin, Update, PEM, ADC common mode rejection with new seismometer connections setup.pdfseismometerASD.pdf
                                           Reply  Wed May 16 21:02:22 2018, Kevin, Update, PEM, ADC common mode rejection with new seismometer connections ADC_noise.pdf
                                        Reply  Thu Jun 14 15:24:32 2018, Steve, Update, PEM, ADC DAC In Line Test Boards are in ADC_DAC_in_(1).JPG
    Reply  Tue Jan 2 15:43:35 2018, Steve, Update, VAC, pumpdown after power outage pumpdown_from_44_mTorr.png
       Reply  Wed Feb 7 15:50:42 2018, Steve, Update, VAC, IFO pressure monitoring CC_analog.png
Message ID: 13755     Entry time: Mon Apr 16 22:09:53 2018     In reply to: 13510     Reply to this: 13756  
Author: Kevin 
Type: Update 
Category: General 
Subject: power outage - BLRM recovery 

I've been looking into recovering the seismic BLRMs for the BS Trillium seismometer. It looks like the problem is probably in the anti-aliasing board. There's some heavy stuff sitting on top of it in the rack, so I'll take a look at it later when someone can give me a hand getting it out.

In detail, after verifying that there are signals coming directly out of the seismometer, I tried to inject a signal into the AA board and see it appear in one of the seismometer channels.

  1. I looked specifically at C1:PEM-SEIS_BS_Z_IN1 (Ch9), C1:PEM-SEIS_BS_X_IN1 (Ch7), and C1:PEM-ACC_MC2_Y_IN1 (Ch27). All of these channels have between 2000--3000 cts.
  2. I tried injecting a 200 mVpp signal at 1.7862 Hz into each of these channels, but the the output did not change.
  3. All channels have 0 cts when the power to the AA board is off.
  4. I then tried to inject the same signal into the AA board and see it at the output. The setup is shown in the first attachment. The second BNC coming out of the function generator is going to one of the AA board inputs; the 32 pin cable is coming directly from the output. All channels give 4.6 V when when the board is powered on regardless of wheter any signal is being injected.
  5. To verify that the AA board is likely the culprit, I also injected the same signals directly into the ADC. The setup is shown in the second attachment. The 32 pin cable is going directly to the ADC. When injecting the same signals into the appropriate channels the above channels show between 200--300 cts, and 0 cts when no signal is injected.
Attachment 1: AA.jpg  175 kB  Uploaded Mon Apr 16 23:33:21 2018  | Hide | Hide all
AA.jpg
Attachment 2: ADC.jpg  161 kB  Uploaded Mon Apr 16 23:37:08 2018  | Hide | Hide all
ADC.jpg
ELOG V3.1.3-